# Channel Choice and Customer Value

### v1

@@ F.4.1 Appendix to Covid Analysis - Re-defining the Organic Adopter Cohort
F.4.2 Appendix to Covid Analysis - Re-defining the Covid Adopter Cohort
In this section, we provide an alternative definition of covid adopters cohort for the covid analysis. We provide a robustness check by redefining covid adopters as those customers who were offline-only before 2019-12-31 and made their first online purchase between 2020-05-01 and 2020-06-30 (i.e., two months later than the exact onset of COVID-19). The definition of organic adopters remains the same as the definition in §5.1.1. Under this definition, we continue to have 12,799 organic adopters but now have 27,448 covid adopters. As before, we exclude the periods used for defining these two cohorts (i.e., 2020-01-01 to 2020-06-30), re-run the same set of DID models as in the main analysis, and present the results in Table A51. Table A52 and A53 present DID results for other total and offline metrics. We find that all results are consistent with the main analysis.
@@ F.4.2 Appendix to Covid Analysis - Re-defining the Covid Adopter Cohort
F.4.3 Appendix to Covid Analysis - Comparing Covid Adopters (March/April 2020) vs. Later Covid Adopters (May/June 2020)
Table A54 compares covid adopters (i.e., adopt during March/April 2020) and later covid adopters (i.e., adopt during May/June 2020). Although the later covid adopters spend more than the covid adopters by 9.96 MCU, it is not large in magnitude and with a p-value of 0.048. So we conclude there is no major difference in post-adoption spend, share of offline spend (i.e., channel utilization), and profit.
@@
F.5 Appendix to Using Natural Log Scale Dependent Variables
@@
Within R2 $2.16\times 10^{-7}$ $4.09\times 10^{-6}$ $6.85\times 10^{-6}$ $5.62\times 10^{-6}$ $7.22\times 10^{-6}$ $9.86\times 10^{-6}$ $9.94\times 10^{-6}$
@@
Table A56 and A57 provide DID results between covid adopters and organic adopters with natural log scale dependent variables for total and offline metrics, respectively.
@@
Table A58 and A59 provide DID results between Black Friday and organic adopters with natural log scale dependent variables for total and offline metrics, respectively.
@@
Table A60 and A61 provide DID results between loyalty program and organic adopters with natural log scale dependent variables for total and offline metrics, respectively.
@@
F.6 Appendix to Handling Zero-Inflated Dependent Variables

### v2

@@ E.5.1 Appendix to Covid Analysis - Re-defining the Organic Adopter Cohort
E.5.2 Appendix to Covid Analysis - Re-defining the Covid Adopter Cohort
In this section, we provide an alternative definition of covid adopters cohort for the covid analysis. We provide a robustness check by redefining covid adopters as those customers who were offline-only before 2019-12-31 and made their first online purchase between 2020-05-01 and 2020-06-30 (i.e., two months later than the exact onset of COVID-19). The definition of organic adopters remains the same as the definition in §5.2.1. Under this definition, we continue to have 12,799 organic adopters but now have 27,448 covid adopters. As before, we exclude the periods used for defining these two cohorts (i.e., 2020-01-01 to 2020-06-30), re-run the same set of DID models as in the main analysis, and present the results in Table A45. Table A46 and A47 present DID results for other total and offline metrics. We find that all results are consistent with the main analysis.
@@ E.5.2 Appendix to Covid Analysis - Re-defining the Covid Adopter Cohort
E.5.3 Appendix to Covid Analysis - Comparing Covid Adopters (March/April 2020) vs. Later Covid Adopters (May/June 2020)
Table A48 compares covid adopters (i.e., adopt during March/April 2020) and later covid adopters (i.e., adopt during May/June 2020). Although the later covid adopters spend more than the covid adopters by 9.96 MCU, it is not large in magnitude with a p-value of 0.048. So we conclude there is no major difference in post-adoption spend, share of offline spend (i.e., channel utilization), and profit.
@@
E.6 Appendix to Using Natural Log Scale Dependent Variables
@@
Within R2 $2.16\times 10^{-7}$ $4.09\times 10^{-6}$ $6.85\times 10^{-6}$ $5.62\times 10^{-6}$ $7.22\times 10^{-6}$ $86\times 10^{-6} 9.9$$94\times 10^{-6}$
@@
Table A50 and A51 provide DID results between covid adopters and organic adopters with natural log scale dependent variables for total and offline metrics, respectively.
@@
Table A52 and A53 provide DID results between Black Friday and organic adopters with natural log scale dependent variables for total and offline metrics, respectively.
@@
Table A54 and A55 provide DID results between loyalty program and organic adopters with natural log scale dependent variables for total and offline metrics, respectively.
@@
E.7 Appendix to Handling Zero-Inflated Dependent Variables